### v1

# A Neuroscience-Inspired Dual-Process Model of Compositional Generalization

Abstract
[…]
footnotetext: Equal contribution.footnotetext: Equal advising.
1 Introduction
Humans can systematically generalize to novel situations by composing familiar concepts in new ways Lake et al. (2017); Spelke and Kinzler (2007); Fodor (1975). This compositional flexibility is most evident in natural language, but is present more broadly in abstract reasoning and even everyday thought, beginning early in life Pomiechowska et al. (2024); Piantadosi et al. (2018); Dautriche and Chemla (2025). As a concrete example, knowing how to “jump”, “turn left”, and do something “twice” allows us to understand “jump twice” or “turn left after jumping.” This capacity for compositional generalization is a hallmark of human intelligence but has proven difficult to replicate in artificial systems Fodor and Pylyshyn (1988); Lake and Baroni (2018). While large language models (LLMs) have achieved remarkable success on various NLP tasks, their ability to systematically generalize, especially on tasks requiring strong compositional reasoning, remains limited Xu et al. (2024); Wang et al. (2024).
Survival in complex environments requires precisely these complementary cognitive abilities: the capacity to rapidly recognize familiar patterns and the ability to flexibly recombine known concepts into novel configurations. The mammalian brain achieves this balance through specialized neural systems that have evolved to handle different aspects of information processing. The hippocampus rapidly encodes episodic experiences, while the neocortex gradually extracts statistical regularities across many encounters, forming what neuroscientists call complementary learning systems (CLS) McClelland et al. (1995); Kumaran et al. (2016). This division enables both quick adaptation to new situations and stable accumulation of generalizable knowledge.
During cognitive processing, these systems engage in a nuanced interplay. The neocortex, basal ganglia, and related structures support intuitive, automatic processing (System 1), while a specialized circuit between the hippocampus (HPC) and prefrontal cortex (PFC) enables deliberate, structured reasoning (System 2) (Figure 1). The HPC-PFC complex is particularly crucial for compositional thinking. The hippocampus rapidly binds episodic elements to form cognitive maps of specific environments Moscovitch et al. (2016); Sun et al. (2025); George et al. (2021); Whittington et al. (2020), while the PFC, through iterative interactions with these episodic representations, extracts regularities across experiences to generate abstract schemas Bein and Niv (2025). This complementary process allows the HPC-PFC circuit to orchestrate systematic application of learned knowledge through step-by-step reasoning Koechlin et al. (2003); Tse et al. (2007); Euston et al. (2012); Preston and Eichenbaum (2013). These schemas, consolidated abstractions derived from experience, provide reusable building blocks that can be recombined to solve novel problems.
Figure 1: Dual-process systems in the brain. Left: Brain regions associated with different cognitive processing systems. The hippocampus (HPC) and prefrontal cortex (PFC) form a complex supporting deliberate reasoning (System 2), while the wider neocortex and related regions support intuitive processing (System 1). Right: Functional characteristics of each system. System 1 (Neocortex & related brain regions) is characterized by fast, intuitive pattern recognition and statistical learning. System 2 (HPC-PFC complex) enables slow, deliberate structured reasoning and schema extraction. These complementary neural systems provide the biological foundation for compositional generalization in human cognition.
Drawing on these neuroscientific insights, we present Mirage (Meta-Inference with Rules and Abstractions from Generalized Experience), a dual-process computational framework designed to achieve robust compositional generalization. Mirage mirrors the brain’s functional organization through two complementary systems:
  1. (1)
    System 1: Meta-Transformer-based Decomposer: A standard Transformer architecture Vaswani et al. (2017) that performs fast, parallel pattern recognition analogous to neocortical processing.
  2. (2)
    System 2: HPC-PFC-inspired Schema Engine: A deliberate reasoning component that extracts reusable schemas, assigns priorities to resolve ambiguities, creates temporary bindings for schema arguments, and executes structured transformations, mirroring HPC-PFC functionality.
These systems operate in coordinated iteration: System 2 identifies relevant schemas and their priorities, injects this structured information into System 1’s context window, and the Transformer uses this schema-augmented input to decompose complex problems into manageable subproblems. After each processing step by System 1, System 2 manages schema application and placeholder substitution, refining the representation for subsequent Transformer processing. This iterative schema-guided refinement allows Mirage to systematically tackle compositional tasks that would otherwise remain intractable.
We evaluate Mirage on the SCAN benchmark Lake and Baroni (2018), which is specifically designed to test compositional generalization. Our key contributions are: (1) The Mirage framework, which integrates a fast, intuitive Transformer-based processor with a deliberate, rule-based schema mechanism in a neuroscience-inspired architecture; (2) Demonstration of strong compositional generalization on SCAN, achieving $>$99% accuracy across all splits; (3) Ablation studies confirming that Mirage’s compositional capabilities critically depend on schema priority management, iterative refinement, and the quality of extracted schemas.
Figure 2: Mirage architecture showing the interplay between its two complementary systems inspired by neuroscience: (A) System 1 (Neural Decomposer) trained on diverse artificial grammars implements neocortex-like pattern recognition, while System 2 (Schema Engine) models HPC-PFC functions by extracting, storing, and applying schemas with priorities. (B) The information processing flow during inference, where the meta-trained transformer performs iterative decomposition of complex expressions (e.g., “walk right twice after turn left"), guided by the schema library. With each pass, System 1 processes one layer of the composition tree and stores bindings in episodic memory, while System 2 handles schema selection and expansion. This dual-process design enables Mirage to systematically solve compositional problems through rule-guided hierarchical processing.
2 Methodology
Intuition.
Mirage consists of two complementary components, an explicit, rule-based Schema Manager (System 2) and a meta-learned Transformer Decomposer (System 1), in a tight feedback loop that mirrors the HPC-PFC division of labor: rapid episodic binding versus deliberate schema-level manipulation. Figure 2 sketches the pipeline; below we formalize its elements and interactions.
2.1 Problem setting
We consider supervised sequence-to-sequence learning from an input space $\mathcal{X}$ to an output space $\mathcal{Y}$, where both are sequences of tokens. Each training example $(x,y)\!\in\!\mathcal{X}\!\times\!\mathcal{Y}$ is generated by a term rewriting system $\mathcal{G}$ (a “grammar”). Term rewriting systems consist of rewrite rules $\mathbf{x}\!\mapsto\!\mathbf{y}$ specifying that the string $\mathbf{x}$ rewrites into $\mathbf{y}$, where each string can contain variables. When such rules are allowed to apply cyclically to their own outputs, this formalism becomes Turing-complete, making it a powerful and generic starting point. We can fully specify such term rewriting system via a finite set of schemas. A schema is an $r$-ary function
$\sigma_{M}:(t_{1},\dots,t_{r})\;\longmapsto\;s_{1}\dots\ s_{q}\ M\,s_{q+1}\,\dots\,s_{r}$
where every placeholder or argument $t_{i}$ is bound at application time to a token or sub-schema $s_{j}$ in composition. A schema is triggered in a sequence by a corresponding modifier token, $M$. For example, in the SCAN grammar, $\sigma_{\textsc{after}}(a,b)=a\;\textit{after}\;b$, where after represents the modifier token. A schema may be composed of multiple sub-schemas, too, becoming a compound schema. For example, we might specify $\sigma_{\textsc{after-twice}}(a,b,c)=a\;\textit{after}\;b\;\textit{twice}$. A grammar may contain compound schemas, but must contain a standard atomic schema for every recognized modifier token.
To preserve deterministic compositional generation and resolvement, we must order the schemas in a grammar by tagging them with an integer priority $\pi(\sigma)$ (e.g. “turn left” must be executed before “twice”). This necessity may be best explained through the simple analogy to ‘PEMDAS’ order of operations in mathematics. Specific operations must take priority to ensure consistent solutions.
Lastly, a grammar must contain an evaluator which executes some function on its arguments. While a schema outputs a sequence, an evaluator $\psi$ maps a schema and arguments to an output sequence that may or may not consist of arguments to the schema.
$\psi:(\sigma,(t_{1},\dots,t_{r}))\;\longmapsto\;e_{1}\dots\ e_{n}$
Hence a grammar is the 3-tuple $\mathcal{G}=(\Sigma,\pi,\psi)$ with $\Sigma$ the schema set. Through recursion, a grammar can output sequences of arbitrary length.
2.2 System 1: Meta-Learned Transformer Decomposer
System 1 is a decoder-only Transformer $\mathcal{T}_{\theta}$ that does a single step of compositional reasoning, and which we apply cyclically to its own outputs to solve reasoning problems. Specifically, it is trained to perform single-step decomposition of any input sequence generated by an arbitrary grammar $\mathcal{G}$, given the definition of $\mathcal{G}$ in-context. This allows compositional problems to be reasoned about in sequential steps, with more complex (i.e. deeper) compositional problems requiring more steps. At inference time, this results in repeated applications of the System 1 transformer to fully decompose a compositional problem and return a result.
Breaking a compositional reasoning task into steps in this manner allows small models to execute complex compositional tasks without explicitly modeling compositional logic directly. Instead, models must only execute ‘prioritized pattern matching,’ or the identification of schema input patterns according to schema priority. This problem is more tractable than modeling k-deep composition, yet is sufficient to achieve systematic compositionality when used iteratively.
2.3 Training Algorithm: Meta-Learning Generalizable Single-Step Decomposition
Sequences with composition can be represented as composition trees. Every non-leaf node in a composition tree built from a sequence must refer to a modifier or action and every leaf node must refer to a primitive. Each SCAN sequence represents a single composition tree. Single-step decomposition for a compositional sequence is equivalent to reducing the depth-level of the composition tree by 1 after resolving all sub-trees that extend into that deepest level.
To parameterize a general function for single-step decomposition of any composition tree defined by a valid grammar, we meta-learn $\mathcal{T}_{\theta}$ on an endless stream of randomly generated grammars $\{\mathcal{G}_{m}\}_{m=1}^{\infty}$ (cf. Lake and Baroni (2018, 2023); Appendix A.1A.2). Training samples to the model consist of the ordered concatenation of the following 3 components: 1) A token sequence representation for a grammar as a set of schemas with corresponding priorities, 2) a maximum 2-deep composition input sequence of modifiers and atomic primitive tokens, generated by the grammar, and 3) the input sequence with single-step decomposition applied, denoted by replacing modifiers with schema names.
$\underbrace{\text{schema-library}}_{\texttt{Schema 1}\,\cdots\texttt{Schema M}}\;\texttt{LP\_SEP}\;\underbrace{\text{input sequence}}_{\smash{\textstyle x}}\;\texttt{SEP}\;\underbrace{\text{target sequence}}_{\smash{\text{DECOMPOSE}(x)}}\;\texttt{EOS}$
2.4 Inference Algorithm: Iterative Refinement with Replacement
Given a trained $\mathcal{T}_{\theta}$, we can reconstruct output sequences from input sequences of any composition depth and generated by any arbitrary grammar. In summary, we take input sequence $x^{(k)}$ with $k$ levels of schema composition and apply our model $\mathcal{T}_{\theta}$. Given the model output, we search for ‘schema name tokens’, which the transformer outputs to represent an instance of decomposition being applied. For every schema name token, we replace that token, as well as all of the arguments required for the schema referenced by that name, with a single atomic primitive. We store this replacement in a ‘replacement map’ (simple hash map or table). This results in a new sequence, $x^{(k-1)}$, with $k-1$ levels of composition (the last layer or bottom layer of the composition tree has been removed). We then apply the model to the new sequence after replacement and decompose the next layer. This zero-shot-given-a-grammar inference procedure is fully defined by Algorithm 1.
Algorithm 1 Zero-Shot Inference on Single Input Sequence Example
1:Trained transformer $\mathcal{T}_{\theta}$, grammar $\mathcal{G}$, input sequence $x^{(k)}$ with composition depth $k$
2:Final flat output sequence $\mathcal{S}_{\mathrm{final}}$ and replacement map $\lambda$
[…]
5:for $d=k,k-1,\dots,1$ do
6:  $y\leftarrow\mathcal{T}_{\theta}\bigl{(}x^{(d)}\bigr{)}$triangleright predict next-level decomposition
7:  while there is a schema token $s$ in $y$ do
8:   let $(\sigma_{s},\,a_{1},\dots,a_{m})$ be the schema name and its $m$ argument tokens in $y$
9:   $p\leftarrow\texttt{ApplySchema}(\sigma_{s},\,a_{1},\dots,a_{m})$triangleright collapse into one primitive
10:   $\lambda\bigl{[}\sigma_{s}(a_{1},\dots,a_{m})\bigr{]}\leftarrow p$
11:   replace the subsequence $\langle\sigma_{s},a_{1},\dots,a_{m}\rangle$ in $y$ with $p$
12:  $x^{(d-1)}\leftarrow y$triangleright one less level of composition
13:return $\mathcal{S}_{\mathrm{final}}\leftarrow x^{(0)},\;\lambda$
After receiving the output from algorithm 1, we simply resolve or unwind the complete value for $S_{final}$, replacing any tokens in the sequence that match keys in $\lambda$ with their corresponding value in $\lambda$.
The number of model iterations required to completely resolve a schema is equal to the depth of the composition tree, where the root node sits at depth 0. Consider an input sequence of length $N$, with a max number of arguments per schema of $A$ (in SCAN, $A=2$). Then, the depth of the tree is $O(\log_{A}N)$, and as such, the number of decomposition or reasoning steps is generally logarithmic in sequence length. This allows processing extremely large compositional sequences in a small number of reasoning steps.
2.5 System 2 details: complementary schema extractors
Schema extraction is treated as a modular component whose internal strategy can be varied without altering the rest of Mirage. We currently provide two contrasting implementations: a symbolic search procedure and a learnable graph-based model built atop a recent computational model of the hippocampus George et al. (2021). Their different strategies make it unlikely that both will fail on the same input, and either one can be replaced as improved algorithms or neuroscientific insights become available.
Extractor Option One: CSCG-inspired extractor
Following the neuroscientific inspiration of the two-component model, we utilize Clone-Structured Causal Graph (CSCG) George et al. (2021), effective in modeling cognitive maps in mice Sun et al. (2025). Adapted from Cloned Hidden Markov Models (CHMM) Dedieu et al. (2019), CSCGs efficiently represent complex sequences via a discrete latent state model. CSCGs further constrain each latent state to deterministically emit a single token, effectively ‘cloning’ each token into a large number of hidden states. Critically, CSCGs can rebind these emission matrices to generalize to patterns analogous to those they were trained on, allowing them to learn simple algorithms Swaminathan et al. (2023). Our approach works by extending this rebinding capability.
The original rebinding mechanism triggers upon encountering unexpected tokens and depends on stable “anchor” tokens that are rarely rebound. We extend this mechanism to bidirectional contexts where novel primitives might precede anchors. Thus, we introduce an adapted rebinding algorithm to explicitly detect anchors and rebinding slots within such contexts, defined formally below.
[…]
1. Episode decoding. We train a CHMM on the raw token stream. Viterbi decoding yields a latent state trace for each demonstration, forming episodes $e=\langle\mathbf{x},\mathbf{y},\mathbf{z}\rangle$ where $\mathbf{z}$ is the state sequence. These episodes serve as structured surrogates of hippocampal encodings.
2. Schema discovery and pruning. Candidate schemas are generated by aligning episode pairs $(e_{i},e_{j})$ using the longest common subsequence (LCS) under a constraint enforcing consistent variable bindings. A candidate is accepted if it validates on all supporting episodes via forward simulation and is supported by at least min_support examples, filtering out ambiguous or redundant rules. Surviving schemas are deduplicated using a keyword–coverage heuristic and pruned iteratively if compositionally subsumed by higher-support schemas. We visualize the set of extracted schemas in Appendix Figure 4.
Algorithm 2 Extraction of compositional schemas
1: Demonstrations $\mathcal{S}$, minimum support $k$
2:Train CHMM on $\mathcal{S}$; decode each $s\in\mathcal{S}$ to obtain episodes $\mathcal{E}$
[…]
4:for all $(e_{i},e_{j})\in\binom{\mathcal{E}}{2}$ do
5:  $\tau\leftarrow\textsc{Align}(e_{i},e_{j})$triangleright injective LCS with variables
6:  if $\tau\not=\bot$ then
7:   $\text{supp}\leftarrow\{e\in\mathcal{E}\mid\textsc{Validate}(\tau,e)\}$
8:   if $|\text{supp}|\geq k$ then $\mathcal{C}\leftarrow\mathcal{C}\cup\{\langle\tau,\text{supp}\rangle\}$
[…]
11:return $\mathcal{S}_{\text{final}}$
Schemas are applied via a greedy, pass-based interpreter that replaces the longest matching input spans until convergence. A derivation is successful if all placeholders are resolved. In compositional cases like SCAN, however, overlapping operator-style schemas (e.g., turn right thrice) may conflict. To resolve this, we induce a precedence relation $\succ$ over $\mathcal{S}_{\text{final}}$.
Inspired by CSCG-style rebinding, we use comparative demonstrations split by <SEP> and utilize one-step resolution. If schema $s_{i}$ fires before $s_{j}$, we record $s_{i}\succ s_{j}$. Aggregating such wins produces a matrix $C\in\mathbb{N}^{|\mathcal{S}_{\text{final}}|\times|\mathcal{S}_{\text{final}}|}$, whose Copeland score
$\text{score}(s_{i})=\sum\nolimits_{j\neq i}\bigl{[}\![\,C_{ij}>0\,]\!-\![\,C_{ji}>0\,]\bigr{]}$
defines a total preorder. Full details are in Appendix B.1.
With precedence learning in place our system (i) discovers variable-binding rewrite rules, (ii) prunes them for compositional minimality, and (iii) orders overlapping operator schemas so that the following Transformer model can utilize deterministic schema information on previously unseen commands.
Extractor option two: enumerative rule miner.
To complement the CSCG path we include a lightweight symbolic “rule-miner” inspired by enumerative program-synthesis methods Udupa et al. (2013). Starting from the demonstrations alone, it treats each input–output pair as an episodic memory and enumerates simple string-rewrite templates (span-to-token replacements, span re-ordering, and wrapper insertions). Templates are accepted only if, when added to the current library, they repair all of their matches and raise corpus-level exact accuracy; accepted rules are then replayed to expose new residual errors and the loop repeats until coverage stabilizes. The miner also records “fires-before” relations between overlapping rules and returns a topologically-sorted precedence schedule together with the final library of primitives, modifiers, and schemas. The resulting grammar is directly fed into the Schema Engine and gives performance indistinguishable from the CSCG variant, underscoring System 2’s modularity.
3 Results
3.1 MIRAGE Performance on SCAN vs. Baselines
We evaluate Mirage and baselines on the canonical SCAN benchmark, following the exact data splits of Lake and Baroni (2018). Unless stated otherwise, all numbers are means $\pm$ SEM over 4 independent runs. Training, hardware, and hyper-parameter details mirror those in §2.
Table 1 summarizes our main results. Mirage achieves 99.59 $\pm$ 0.24% accuracy on the full task while trained in a task-agnostic manner on randomly generated grammars rather than SCAN-specific examples, and reaches near-perfect performance on every split. Perhaps even more importantly, a single trained Mirage is generalizable to any compositional grammar (given its specification) with appropriate vocabulary size for the trained model (number of modifier tokens, primitives, and arguments per schema must be less than or equal to the corresponding parameters used to train the model).
Table 1: SCAN accuracy (mean  $\pm$ SEM over 4 runs). † Baseline is re-trained on each split’s train set;
SCAN splits
Model Full Simple Length prim_jump temp_or
[…]
In contrast, transformers trained directly on SCAN solve the normal split effectively, but fail to generalize across length and new templates in SCAN. This demonstrates significant overfitting, failing to learn any generalizable notion of composition or compositional reasoning. The Transformer+SC_Library baseline represents a standard transformer trained to evaluate SCAN, with the addition of the SCAN schema library in-context (i.e. pre-pend the SCAN schema library tokens used for inference evaluation of MIRAGE to the training sequences of the transformer).
We compare against a simple baseline by prepending the schema library directly to the input sequence to test whether the Transformer can interpret it without explicit guidance. Interestingly, as shown in Table 1, the opposite holds true: the additional context reduces performance on token-based splits and leads to only minor gains on the length split. This suggests that the model cannot effectively leverage schema information without a dedicated training objective, and length split improvements appear to merely stem from the increased input size rather than meaningful schema usage.
3.2 Ablation Analysis Study
To identify which design choices are essential for MIRAGE, we ran four ablations: (1) removing priority tokens, (2) disabling iterative refinement and forcing single-pass decomposition, (3) training on sequences with unbounded composition depth, and (4) evaluating with imperfectly extracted grammars. Each change substantially degraded accuracy, underscoring the importance of explicit precedence cues, step-wise decomposition, bounded-depth meta-training, and a correct schema specification. We have the following key observations:
  1. 1.
    Priority Tokens are Critical: Eliminating explicit priority scheduling reduces full-task accuracy from 99.59 ± 0.24% to 71.92 ± 0.72%. Although the model still solves a majority of commands, the 28-point drop reveals that it often chooses the wrong schema when several overlap. In practice, the transformer falls back on brittle positional cues and succeeds only when those cues coincide with the intended hierarchy. Placing schemas in a fixed order without dedicated PRIORITY_ tokens proved insufficient during development, confirming that the network needs explicit embeddings to internalize precedence. Prioritization therefore remains the hardest part of single-step decomposition; giving the model its own priority-token embedding space was crucial for reliable generalization.
  2. 2.
    Iterative Refinement Process Facilitates Generalization: We experimented with parameterizing complete single-shot decomposition prior to switching to single-step decomposition with the iterative refinement algorithm described above. In our experiments, we failed to train transformers to systematically learn or generalize complete single-shot decomposition algorithms even when trained on $k$-deep composition example sequences (2-deep train sequences also failed). The iterative refinement algorithm allows transformer models to learn more generalizable rules focused solely on prioritized schema identification, as opposed to full decomposition. In this sense, the iterative refinement algorithm seems important for facilitating generalized inference.
  3. 3.
    Training with Unlimited Depth Composition Sequences Fails: While we cannot present specific evidence as to why this might be the case, experiments repeatedly revealed that models (especially, larger models) trained on sequences including $k$-deep composition (limited only by sequence context length) tended to perform worse on SCAN, reaching a standard peak performance of less than 50%. Training on 2-deep composition sequences ultimately proved more generalizable even when performing inference on k-deep composition sequences, like in SCAN. This may be the result of models trained on deeper sequences attempting to model composition more closely, which ultimately becomes detrimental, as general patterns are not observed. A more focused learning paradigm on prioritized pattern matching only proves more generalizable and effective.
  4. 4.
    Correct Grammar Libraries are Necessary: Introducing stochastic boundary noise, that is, random shifts of one or two positions in num_args_before or num_args_after for every schema, drops full-task accuracy from 99.59 ± 0.24% to 0.065 ± 0.021%. Across four random seeds, the accuracy never exceeded 0.11 percent. This sharp decline shows that accurate schema extraction is essential; even small mistakes in the grammar cause the model’s compositional reasoning to break down.
3.3 Experimental Takeaways
Our findings support the hypothesis that coupling an explicit schema extraction system with a fast neural processor is necessary and (empirically) sufficient for systematic compositionality. Our contributions may be summarized through the following key points. (1) A HPC-PFC-inspired model consisting of corresponding components can effectively solve compositional generalization tasks zero-shot, when given a complete grammar specification. (2) The HPC-PFC complex, System 2, may be modeled with a wide variety of potential schema extraction methods. We present two methods based on program synthesis and neuro-inspired CSCG models. (3) Meta-learning transformers on an infinite random grammar stream with only 2-level deep composition training examples is sufficient to perform generalized single-step decomposition on arbitrary grammars in k-deep composition scenarios. Via an iterative refinement process, a trained model $\mathcal{T}_{\theta}$ can effectively decompose length $N$ compositional sequences in $O(\log N)$ reasoning steps. (4) We provide ablation results demonstrating the importance of the 2-system design and other components integrated into our core system. We believe that these contributions, specifically the overall framework of our system design and the use of transformers in an iterative reasoning/refine process, present intriguing areas of research moving forward. We believe that developing principled neuro-inspired AI systems represents perhaps the most tangible pursuit of the type of systematic generalization to new contexts and scenarios that humans excel at and even massive scale AI has not yet seemed to achieve.
4 Related Work
Adjacent Work on Compositionality
Several research directions have tackled compositional generation, spanning auxiliary objectives, neuro-symbolic models, prompting strategies, and mechanistic Transformer analyses. Jiang and Bansal (2021) use auxiliary supervision to improve compositional generalization in Transformers, by introducing sequence prediction tasks that encourage structural understanding, finding that less contextualized representations improve SCAN performance. Another strategy, LANE Liu et al. (2020), learns analytical expressions and leverages memory augmentation inspired by variable-slot reasoning from cognitive science. Though related to our extractor component, LANE is non-Transformer-based and follows a distinct training paradigm, yet still achieves strong SCAN results. Neuro-symbolic methods combine neural and symbolic reasoning for compositionality. The Compositional Program Generator (CPG) Klinger et al. (2024) integrates grammar-based modularity and abstraction, aligning closely with Mirage’s schema-based extraction strategy. Similarly, the Neural-Symbolic Recursive Machine (NSR) Li et al. (2024) learns syntax and semantics jointly via a Grounded Symbol System, using tree structures reminiscent of our Transformer-based neural processor.
Compositionality in Deep Learning Structures
Finally, understanding the capabilities and limitations of existing architectures, particularly Transformers, is essential for developing more effective models.  Russin et al. (2024) survey recent work on compositionality in deep neural networks (DNNs), exploring how DNNs, particularly large language models (LLMs), can achieve compositional generalization to a specific degree through architectural inductive biases and metalearning.  Wang et al. (2024) investigate Transformers’ ability to perform implicit reasoning, finding that they can learn these skills only through "grokking", but struggle with systematic generalization in compositional tasks.  Brinkmann et al. (2024) perform a mechanistic analysis of a Transformer trained on a synthetic reasoning task, identifying interpretable mechanisms like backward chaining. Overall, these works highlight the challenges in achieving robust compositional reasoning with standard transformers, further motivating the need for specialized architectures like Mirage.
Neuroscience Inspiration
We further take some to reflect upon work in neuroscience and cognition to discuss the inspiration Mirage was built upon.  Kurth-Nelson et al. (2022) propose that hippocampal replay implements compositional computation by assembling entities into relationally-bound structures, aligning with Mirage’s schema application and transformation mechanisms. Their hypothesis strengthens the neuroscientific grounding of Mirage by linking it to the compositional role of replay in the brain. The HPC-PFC circuits, conserved across mammalian evolution Murray et al. (2017), represent nature’s time-tested solution to compositional reasoning, offering a bridge between neuroscience and AI that may illuminate new paths toward more human-like intelligence.
5 Conclusion
We introduced Mirage, a neuroscience-inspired dual-process framework that explicitly separates fast pattern matching from deliberate schema manipulation. Meta-learning a single-step Transformer decomposer over a stream of random grammars, combined with a priority-aware schema manager, yields state-of-the-art zero-shot compositional generalization, given a complete grammar: Mirage attains 99.6% accuracy on the full SCAN task and performs similarly on the other splits, without exposing SCAN training data to the Transformer.
Beyond outperforming strong baselines, our ablation study confirms that (i) priority scheduling and (ii) iterative refinement are each critical: Removing either component collapses systematicity.
Looking more broadly at the reasoning landscape today, dominant approaches operate by training large homogeneous neural networks to perform step-by-step inference Guo et al. (2025); Muennighoff et al. (2025); OpenAI et al. (2024). Our model should be seen as an alternative route. We learn a small step-by-step reasoner, and augment it with an external module inspired by the modular architecture of the brain. Our model also sits within the paradigm of separating knowledge from inference, which is a staple of classic logical reasoning. Looking forward we hope that the judicious application of modularity, particularly inspired by brain architecture, could prove a promising complementary route for learning to reason.
Limitations. We initially consider a radically general knowledge representation—term rewriting systems—but then quickly restrict it to make the model tractable. We are optimistic however that taking Turing-complete representations as a starting point could prove valuable in the future. As our modeling goals are primarily inspired by neuroscience, we also take a noncommittal stance regarding the optimal algorithm for learning or extracting schemas, and provide two distinct algorithms. For CSCG extraction we further assume existence of atomic demonstrations. Last, we only evaluate on SCAN, because it is a canonical check of compositionalty Lake and Baroni (2023), but we hope that by moving toward increasingly expressive computational formalisms for schemas, our architecture could find further application.
Future Directions and Broader Impacts.
We plan to (i) embed Mirage as a plug-in reasoner/planner for LLMs, while aiming to enhance their abilities to more robustly build and use world models; and (ii) examine its internal states to test neuroscientific hypotheses about HPC–PFC interactions. Broader impact: by exposing an inspectable schema layer, Mirage can make neural reasoning more transparent and auditable, however, deployments, particularly in safety-critical settings, should include clear usage policies and routine monitoring.
Appendix
Appendix A Methodology details
A.1 Transformer Model Vocabulary Specifications for Meta-Learning
The vocabulary of our transformer $\mathcal{T}_{\theta}$ is defined with the following components. To use the model with any arbitrary grammar with its own specific vocabulary, like SCAN, tokens must be anonymized to match this format. For example, in SCAN, ‘primitives’ are tokens like ‘walk’, ‘jump’, ‘look’, etc., while modifiers are tokens ‘and’, ‘after’, ‘opposite left’, or ‘twice’. Anonymizing a grammar in this way is a simple, deterministic process.
  1. 1.
    Primitives: PRIM_0, …, PRIM_P. Primitives are the basic atomic elements in a vocabulary. Integer parameter $P$ sets the number of allowed unique primitives.
  2. 2.
    Modifiers: MOD_0, …, MOD_M. Schemas bind to and define the action of modifier tokens as functions. A given grammar $\mathcal{G}$ defines the action of a modifier in a sequence generated under it. Integer parameter $M$ sets the number of allowed unique modifiers.
  3. 3.
    Argument Tokens: ARG_0, …, ARG_2*A. Argument tokens are used in the in-context token sequences used to represent grammars. Specifically, argument tokens define the placeholders that a schema could bind to. For example, a schema may be represented presented via a token sequence as ARG_0, …, ARG_A MOD_0 ARG_A+1, …, ARG_2A, where the argument tokens are placeholders for real arguments to the schema and the modifier token represents the action the schema is binding to. Integer parameter $A$ sets the maximum number of arguments a schema can have that occur before and after its principle modifier.
  4. 4.
    Schema Name Tokens: SC_0, …, SC_S. These tokens define the names of specific schemas in the grammar or $|\Sigma|$. These names occur in-context to serve as markers for bindings in output sequences after a model application. Integer parameter $S$ controls the number of schemas in a grammar, or $|\Sigma|$. We must have $S\geq m$ for a grammar G, as each modifier must occur in at least one schema to be well-defined.
  5. 5.
    Priority Tokens: PRIORITY_0, …, PRIORITY_S. These tokens define the priority with which schemas are to be bound to. These tokens occur next to schema name tokens in-context to define their priority.
  6. 6.
    Administrative Tokens: EOS, SEP, SC_DEF, SC_PRI, SC_SEP, LP_SEP , PAD. These tokens are for formatting grammars in-context with input sequences generated by them.
A.2 Training Algorithms
During the transformer training process, new random grammars are generated at regular intervals. These grammars are added to a ‘Grammar Buffer’ maintaining the full set of previously generated grammars during the training process. At every step, the transformer samples grammars from this buffer and delegates the construction of random 2-deep composition input sequences and their corresponding output sequences. The concatenation of these components is then fed into the model. Multiple grammars are generally represented in each batch.
In the Methods section, we referenced a variety of small sub-algorithms that the transformer must delegate during the meta-learning training process. These include actually generating random grammars to add to the buffer at regular intervals (defined by a hyperparameter), generating input/output sequence pairs given a grammar, and actually sampling from the grammar buffer for each batch. Each of these algorithms are detailed here.
  • Random Grammar Generation: For some subset of modifier tokens $M_{G}\subseteq M$, we generate a schema, $\sigma$. For each,
    • Choose $A_{before},A_{after}\leq A_{max}$ for number of arguments before and after the modifier token in the schema output format.
    • Define $\psi(\sigma,(t_{1},\dots,t_{A_{before}+A_{after}}))$.
    • Define $\pi(\sigma)$.
  • Input Sequence Generation: To generate an input sequence from a grammar, choose a schema, $\sigma$. For each argument in $\sigma$, select an atomic primitive or another schema $\sigma_{sub}$ with $\pi(\sigma)<\pi(\sigma_{sub})$. For $\sigma_{sub}$’s, select all primitives as arguments. This generates 2-deep schema composition sequences.
  • Output Sequence Generation: For the deepest layer of composition, simply replace the modifier token $M$ of applied schema $\sigma_{M}$ with a ’schema name token’ like SC_M to represent $\sigma_{M}$. We train our model to output this format so that we can easily detect where the model has performed decomposition. This facilitates the zero-shot inference procedure, given a new grammar specification, outlined below.
  • Grammar Buffer Sampling: A Grammar Buffer consists of a set of $N$ previously used grammars. The buffer maintains the content of these grammars, as well as a corresponding $C_{i}$ representing the number of times a grammar $G_{i}$ has been used. When sampling from the buffer, we sample from the inverse $C_{i}$’s for all buffers. Let $C^{\prime}_{i}$ be $\frac{1}{C_{i}+s}$, for smoothing factor $s,0<s\leq 1$. Then, the probability of selecting $G_{i}$ is $\frac{C^{\prime}_{i}}{\sum_{j=1}^{N}C^{\prime}_{j}}$.
Appendix B Additional details about schema extractors
B.1 Priority Scoring Algorithm for CSCG-extractor
Algorithm 3 Learning operator precedence from one-step demonstrations
[…]
3:for all episode $e\in\mathcal{E}$ do
4:  $(\mathbf{x},\mathbf{y})\leftarrow$ split $e$ at <SEP>
5:  $P\leftarrow\{\,s\in\mathcal{S}\mid\text{pattern}(s)\subset\mathbf{x}\,\}$triangleright schemas present in the input command
6:  $F\leftarrow\{\,s\in P\mid\text{apply}(s,\mathbf{x})=\mathbf{y}\,\}$triangleright schemas that fired in the first step
7:  for all $s_{w}\in F$ do
8:   for all $s_{\ell}\in P\setminus F$ do
9:     $C[s_{w},s_{\ell}]\leftarrow C[s_{w},s_{\ell}]+1$
10:for all $s_{i}\in\mathcal{S}$ do
11:  $\text{score}(s_{i})\leftarrow\sum_{j\neq i}\bigl{(}[\![C_{i,j}>0]\!]-[\![C_{j,i}>0]\!]\bigr{)}$triangleright Copeland score
12:return $\{\text{score}(s_{i})\}_{s_{i}\in\mathcal{S}}$
B.2 Enumerative Rule-Miner Details
This appendix expands on the enumerative rule miner introduced in §2.5.
B.2.1 Goal
From a support set $\mathcal{D}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N}$ of input–output strings, the miner induces a rewrite grammar $G=(\mathcal{P},\mathcal{M},\Sigma,\pi)$: primitives $\mathcal{P}$, modifiers $\mathcal{M}$, schemas $\Sigma=\{\sigma_{1},\dots,\sigma_{S}\}$, and a precedence map $\pi$. Each schema is a variable-binding template whose left-hand side (LHS) may contain string literals, span variables $x_{k}$ (arbitrary substrings), or token variables $u_{k}$ (single words).
B.2.2 Algorithm
  1. (1)
[…]
  2. (2)
    Repair test Insert a candidate template $\tau$ into the current grammar and re-evaluate all demonstrations. $\tau$ is accepted iff it rewrites every match consistently and increases corpus-level exact accuracy; accepted templates are appended to $\Sigma$.
  3. (3)
    Precedence induction During replay, whenever two schemas match the same string and $\sigma_{i}$ fires before $\sigma_{j}$, record $\sigma_{i}\!\succ\!\sigma_{j}$. A topological sort of this graph yields $\pi$.
  4. (4)
    Termination Iterate steps (1–3) until the candidate queue empties, accuracy plateaus, or a fixed budget is reached.
Any alternative proposal mechanism (e.g., beam search, neural scorers, constraint solvers) can replace the enumerator in step (1) without changing the downstream interface: the Schema Engine consumes only $(\mathcal{P},\mathcal{M},\Sigma,\pi)$.
Appendix C CSCG application on SCAN
Figure 3: We apply a Clone-Structured Causal Graph with 100 clones directly on a concatenated subset of SCAN sequences and visualize the resulting model as a simple directed graph.
In line with showcasing the inability of pure Transformers to solve different SCAN splits on its own in section 3, we explored directly applying CSCG on concatenated SCAN sequences. However, due to innate traits of the model architecture, SCAN, or other compositional tasks, quickly exhaust the structural clone bottlenecks. Specifically, the model is unable to distinguish between more than n sequences due to the Separator token, which is contained between every input and output. At the point of entering this token, the model must commit to one specific clone, leading to the loss of prior information. We further experienced with alternative variants that take the state across all clones of a single emission into account, increasing the number of potential configurations that the state can exhibit, but these modifications still failed to meaningfully solve any SCAN split.
Appendix D Compute Resources
Experiments were performed on readily available research hardware: single recent GPUs (e.g., A100 or H100) on a campus cluster and, occasionally, cloud services. Meta-training the Transformer required a few GPU-hours on one card, and each extra seed or ablation consumed a similar budget. Both schema extractors finish in under a minute on a CPU, and evaluating the full SCAN test set completes in under five minutes on a single GPU.
Appendix E CSCG Extractor Schemas
Figure 4: Overview of extracted schemas by the CSCG extractor, visualized as directed sequence graphs. Note that turn is a special case, which evokes different behavior when combined with around or opposite.
The CSCG extractor produces both, a textual representation of the extracted schemas, as well as a graph-based visualization, showcasing the direct correspondence between input and output variables. This visualization approach reveals potential for schema comparison through graph-based representations, as atomic schema demonstrations share consistent structural components. Additionally, the extractor correctly identifies special cases in the "turn" and "around" schemas, recognizing that these produce no additional output when paired with the turn primitive, unlike other primitives such as jump or run.

### v3

@@
 Abstract
-[…]
+Deep learning models struggle with systematic compositional generalization, a hallmark of human cognition.
+We propose Mirage, a neuro-inspired dual-process model that offers a processing account for this ability.
+It combines a fast, intuitive “System 1” (a meta-trained Transformer) with a deliberate, rule-based “System 2” (a Schema Engine), mirroring the brain’s neocortical and hippocampal–prefrontal circuits.
+Trained to perform general, single-step decomposition on a stream of random grammars, Mirage achieves >>99% accuracy on all splits of the SCAN benchmark in a task-agnostic setting.
+Ablations confirm that the model’s systematic behavior emerges from the architectural interplay of its two systems, particularly its use of explicit, prioritized schemas and iterative refinement.
+In line with recent progress on recursive/recurrent Transformer approaches, Mirage preserves an iterative neural update while externalizing declarative control into an interpretable schema module.
+Our work provides a concrete computational model for interpreting how compositional reasoning can arise from a modular cognitive architecture.
 footnotetext: Equal contribution.footnotetext: Equal advising.
 1 Introduction
-Humans can systematically generalize to novel situations by composing familiar concepts in new ways Lake et al.
-(2017); Spelke and Kinzler (2007); Fodor (1975).
-This compositional flexibility is most evident in natural language, but is present more broadly in abstract reasoning and even everyday thought, beginning early in life Pomiechowska et al.
-(2024); Piantadosi et al.
-(2018); Dautriche and Chemla (2025).
-As a concrete example, knowing how to “jump”, “turn left”, and do something “twice” allows us to understand “jump twice” or “turn left after jumping.” This capacity for compositional generalization is a hallmark of human intelligence but has proven difficult to replicate in artificial systems Fodor and Pylyshyn (1988); Lake and Baroni (2018).
-While large language models (LLMs) have achieved remarkable success on various NLP tasks, their ability to systematically generalize, especially on tasks requiring strong compositional reasoning, remains limited Xu et al.
-(2024); Wang et al.
-(2024).
-Survival in complex environments requires precisely these complementary cognitive abilities: the capacity to rapidly recognize familiar patterns and the ability to flexibly recombine known concepts into novel configurations.
-The mammalian brain achieves this balance through specialized neural systems that have evolved to handle different aspects of information processing.
-The hippocampus rapidly encodes episodic experiences, while the neocortex gradually extracts statistical regularities across many encounters, forming what neuroscientists call complementary learning systems (CLS) McClelland et al.
-(1995); Kumaran et al.
-(2016).
-This division enables both quick adaptation to new situations and stable accumulation of generalizable knowledge.
-During cognitive processing, these systems engage in a nuanced interplay.
-The neocortex, basal ganglia, and related structures support intuitive, automatic processing (System 1), while a specialized circuit between the hippocampus (HPC) and prefrontal cortex (PFC) enables deliberate, structured reasoning (System 2) (Figure 1).
-The HPC-PFC complex is particularly crucial for compositional thinking.
-The hippocampus rapidly binds episodic elements to form cognitive maps of specific environments Moscovitch et al.
-(2016); Sun et al.
-(2025); George et al.
-(2021); Whittington et al.
-(2020), while the PFC, through iterative interactions with these episodic representations, extracts regularities across experiences to generate abstract schemas Bein and Niv (2025).
-This complementary process allows the HPC-PFC circuit to orchestrate systematic application of learned knowledge through step-by-step reasoning Koechlin et al.
-(2003); Tse et al.
-(2007); Euston et al.
-(2012); Preston and Eichenbaum (2013).
-These schemas, consolidated abstractions derived from experience, provide reusable building blocks that can be recombined to solve novel problems.
-Figure 1: Dual-process systems in the brain.
-Left: Brain regions associated with different cognitive processing systems.
-The hippocampus (HPC) and prefrontal cortex (PFC) form a complex supporting deliberate reasoning (System 2), while the wider neocortex and related regions support intuitive processing (System 1).
-Right: Functional characteristics of each system.
-System 1 (Neocortex & related brain regions) is characterized by fast, intuitive pattern recognition and statistical learning.
-System 2 (HPC-PFC complex) enables slow, deliberate structured reasoning and schema extraction.
-These complementary neural systems provide the biological foundation for compositional generalization in human cognition.
-Drawing on these neuroscientific insights, we present Mirage (Meta-Inference with Rules and Abstractions from Generalized Experience), a dual-process computational framework designed to achieve robust compositional generalization.
-Mirage mirrors the brain’s functional organization through two complementary systems:
+Humans effortlessly generalize by composing known concepts in new ways, a capacity termed systematic compositionality [1, 2].
+Yet, even large-scale neural networks often fail at such generalization, tending to learn shallow heuristics instead of abstract rules [3, 4, 5].
+This paper provides a processing account for compositional reasoning, inspired by the dual-process architecture of the human brain.
+Cognitive neuroscience suggests that intelligent behavior arises from the interplay of complementary learning systems (CLS) [6, 7].
+This theory posits that a fast, intuitive “System 1,” supported by the neocortex, handles pattern recognition, while a slower, deliberate “System 2,” involving the hippocampus (HPC) and prefrontal cortex (PFC), supports structured, episodic, and goal-directed reasoning [8, 9].
+The HPC rapidly encodes specific experiences (episodes), which the PFC abstracts over time into reusable schemas, generalizable knowledge structures, that enable flexible planning and inference [10, 11].
+This division of labor allows for both rapid adaptation and systematic application of knowledge.
+We instantiate this theory in Mirage (Meta-Inference with Rules and Abstractions from Generalized Experience), a dual-process model with two interacting components:
   1. (1)
-    System 1: Meta-Transformer-based Decomposer: A standard Transformer architecture Vaswani et al. (2017) that performs fast, parallel pattern recognition analogous to neocortical processing.
+    System 1 (Neural Decomposer): A compact Transformer trained via meta-learning to perform a single step of compositional decomposition.
+This parallels the fast, pattern-matching function of the neocortex.
   2. (2)
-    System 2: HPC-PFC-inspired Schema Engine: A deliberate reasoning component that extracts reusable schemas, assigns priorities to resolve ambiguities, creates temporary bindings for schema arguments, and executes structured transformations, mirroring HPC-PFC functionality.
-These systems operate in coordinated iteration: System 2 identifies relevant schemas and their priorities, injects this structured information into System 1’s context window, and the Transformer uses this schema-augmented input to decompose complex problems into manageable subproblems.
-After each processing step by System 1, System 2 manages schema application and placeholder substitution, refining the representation for subsequent Transformer processing.
-This iterative schema-guided refinement allows Mirage to systematically tackle compositional tasks that would otherwise remain intractable.
-We evaluate Mirage on the SCAN benchmark Lake and Baroni (2018), which is specifically designed to test compositional generalization.
-Our key contributions are: (1) The Mirage framework, which integrates a fast, intuitive Transformer-based processor with a deliberate, rule-based schema mechanism in a neuroscience-inspired architecture; (2) Demonstration of strong compositional generalization on SCAN, achieving $>$99% accuracy across all splits; (3) Ablation studies confirming that Mirage’s compositional capabilities critically depend on schema priority management, iterative refinement, and the quality of extracted schemas.
-Figure 2: Mirage architecture showing the interplay between its two complementary systems inspired by neuroscience: (A) System 1 (Neural Decomposer) trained on diverse artificial grammars implements neocortex-like pattern recognition, while System 2 (Schema Engine) models HPC-PFC functions by extracting, storing, and applying schemas with priorities.
-(B) The information processing flow during inference, where the meta-trained transformer performs iterative decomposition of complex expressions (e.g., “walk right twice after turn left"), guided by the schema library.
-With each pass, System 1 processes one layer of the composition tree and stores bindings in episodic memory, while System 2 handles schema selection and expansion.
-This dual-process design enables Mirage to systematically solve compositional problems through rule-guided hierarchical processing.
-2 Methodology
-Intuition.
-Mirage consists of two complementary components, an explicit, rule-based Schema Manager (System 2) and a meta-learned Transformer Decomposer (System 1), in a tight feedback loop that mirrors the HPC-PFC division of labor: rapid episodic binding versus deliberate schema-level manipulation.
-Figure 2 sketches the pipeline; below we formalize its elements and interactions.
-2.1 Problem setting
-We consider supervised sequence-to-sequence learning from an input space $\mathcal{X}$ to an output space $\mathcal{Y}$, where both are sequences of tokens.
-Each training example $(x,y)\!\in\!\mathcal{X}\!\times\!\mathcal{Y}$ is generated by a term rewriting system $\mathcal{G}$ (a “grammar”).
-Term rewriting systems consist of rewrite rules $\mathbf{x}\!\mapsto\!\mathbf{y}$ specifying that the string $\mathbf{x}$ rewrites into $\mathbf{y}$, where each string can contain variables.
-When such rules are allowed to apply cyclically to their own outputs, this formalism becomes Turing-complete, making it a powerful and generic starting point.
-We can fully specify such term rewriting system via a finite set of schemas.
-A schema is an $r$-ary function
-$\sigma_{M}:(t_{1},\dots,t_{r})\;\longmapsto\;s_{1}\dots\ s_{q}\ M\,s_{q+1}\,\dots\,s_{r}$
-where every placeholder or argument $t_{i}$ is bound at application time to a token or sub-schema $s_{j}$ in composition.
-A schema is triggered in a sequence by a corresponding modifier token, $M$.
-For example, in the SCAN grammar, $\sigma_{\textsc{after}}(a,b)=a\;\textit{after}\;b$, where after represents the modifier token.
-A schema may be composed of multiple sub-schemas, too, becoming a compound schema.
-For example, we might specify $\sigma_{\textsc{after-twice}}(a,b,c)=a\;\textit{after}\;b\;\textit{twice}$.
-A grammar may contain compound schemas, but must contain a standard atomic schema for every recognized modifier token.
-To preserve deterministic compositional generation and resolvement, we must order the schemas in a grammar by tagging them with an integer priority $\pi(\sigma)$ (e.g.
-“turn left” must be executed before “twice”).
-This necessity may be best explained through the simple analogy to ‘PEMDAS’ order of operations in mathematics.
-Specific operations must take priority to ensure consistent solutions.
-Lastly, a grammar must contain an evaluator which executes some function on its arguments.
-While a schema outputs a sequence, an evaluator $\psi$ maps a schema and arguments to an output sequence that may or may not consist of arguments to the schema.
-$\psi:(\sigma,(t_{1},\dots,t_{r}))\;\longmapsto\;e_{1}\dots\ e_{n}$
-Hence a grammar is the 3-tuple $\mathcal{G}=(\Sigma,\pi,\psi)$ with $\Sigma$ the schema set. Through recursion, a grammar can output sequences of arbitrary length.
-2.2 System 1: Meta-Learned Transformer Decomposer
-System 1 is a decoder-only Transformer $\mathcal{T}_{\theta}$ that does a single step of compositional reasoning, and which we apply cyclically to its own outputs to solve reasoning problems.
-Specifically, it is trained to perform single-step decomposition of any input sequence generated by an arbitrary grammar $\mathcal{G}$, given the definition of $\mathcal{G}$ in-context.
-This allows compositional problems to be reasoned about in sequential steps, with more complex (i.e.
-deeper) compositional problems requiring more steps.
-At inference time, this results in repeated applications of the System 1 transformer to fully decompose a compositional problem and return a result.
-Breaking a compositional reasoning task into steps in this manner allows small models to execute complex compositional tasks without explicitly modeling compositional logic directly.
-Instead, models must only execute ‘prioritized pattern matching,’ or the identification of schema input patterns according to schema priority.
-This problem is more tractable than modeling k-deep composition, yet is sufficient to achieve systematic compositionality when used iteratively.
-2.3 Training Algorithm: Meta-Learning Generalizable Single-Step Decomposition
-Sequences with composition can be represented as composition trees.
-Every non-leaf node in a composition tree built from a sequence must refer to a modifier or action and every leaf node must refer to a primitive.
-Each SCAN sequence represents a single composition tree.
-Single-step decomposition for a compositional sequence is equivalent to reducing the depth-level of the composition tree by 1 after resolving all sub-trees that extend into that deepest level.
-To parameterize a general function for single-step decomposition of any composition tree defined by a valid grammar, we meta-learn $\mathcal{T}_{\theta}$ on an endless stream of randomly generated grammars $\{\mathcal{G}_{m}\}_{m=1}^{\infty}$ (cf.
-Lake and Baroni (2018, 2023); Appendix A.1A.2).
-Training samples to the model consist of the ordered concatenation of the following 3 components: 1) A token sequence representation for a grammar as a set of schemas with corresponding priorities, 2) a maximum 2-deep composition input sequence of modifiers and atomic primitive tokens, generated by the grammar, and 3) the input sequence with single-step decomposition applied, denoted by replacing modifiers with schema names.
-$\underbrace{\text{schema-library}}_{\texttt{Schema 1}\,\cdots\texttt{Schema M}}\;\texttt{LP\_SEP}\;\underbrace{\text{input sequence}}_{\smash{\textstyle x}}\;\texttt{SEP}\;\underbrace{\text{target sequence}}_{\smash{\text{DECOMPOSE}(x)}}\;\texttt{EOS}$
-2.4 Inference Algorithm: Iterative Refinement with Replacement
-Given a trained $\mathcal{T}_{\theta}$, we can reconstruct output sequences from input sequences of any composition depth and generated by any arbitrary grammar.
-In summary, we take input sequence $x^{(k)}$ with $k$ levels of schema composition and apply our model $\mathcal{T}_{\theta}$.
-Given the model output, we search for ‘schema name tokens’, which the transformer outputs to represent an instance of decomposition being applied.
-For every schema name token, we replace that token, as well as all of the arguments required for the schema referenced by that name, with a single atomic primitive.
-We store this replacement in a ‘replacement map’ (simple hash map or table).
-This results in a new sequence, $x^{(k-1)}$, with $k-1$ levels of composition (the last layer or bottom layer of the composition tree has been removed).
-We then apply the model to the new sequence after replacement and decompose the next layer.
-This zero-shot-given-a-grammar inference procedure is fully defined by Algorithm 1.
+    System 2 (Schema Engine): An explicit symbolic module that extracts, stores, and applies prioritized schemas, managing variable bindings in an episodic memory.
+This mirrors the deliberate, schema-based reasoning of the HPC–PFC loop.
+At inference, these systems engage in an iterative loop: System 2 extracts a compositional grammar from examples and uses it to guide the repeated application of System 1.
+Each pass of the System 1 Transformer refines the problem representation by one level, allowing the model to recursively decompose complex structures.
+This architecture provides an interpretable mechanism for generalization.
+On the SCAN benchmark [4], Mirage achieves state-of-the-art performance across all splits, despite its Transformer component never being trained on SCAN-specific data.
+Our work demonstrates that a modular, neuro-inspired architecture can provide a powerful and interpretable processing account for systematic generalization.
+Figure 1: Mirage architecture and inference loop.
+(A) System 1 (Neural Decomposer) is a meta-trained Transformer for pattern matching.
+System 2 (Schema Engine) extracts and manages a library of prioritized schemas, modeling HPC–PFC function.
+(B) During inference, the systems iterate.
+For a command like “walk right twice after turn left,” System 2 provides the relevant grammar.
+System 1 applies one decomposition step per pass, with outputs stored in an episodic memory.
+This process repeats, reducing the composition tree layer by layer until a primitive action sequence is produced.
+2 A Dual-Process Framework
+Mirage solves compositional tasks through an iterative refinement process managed by its two systems.
+The core idea is to offload declarative knowledge (the "what") to an explicit schema library (System 2) and procedural knowledge (the "how") to a general-purpose neural reasoner (System 1).
+System 2: Schema Engine.
+A compositional task is defined by a grammar $\mathcal{G}=(\Sigma,\pi,\psi)$, where $\Sigma$ is a set of schemas (rewrite rules), $\pi$ is a priority ordering, and $\psi$ is an evaluator for primitive actions.
+System 2’s role is to induce this grammar from a small set of examples.
+This extraction is modular; we show it can be done with either a symbolic rule-miner or a neuro-inspired model based on Clone-Structured Causal Graphs (CSCGs) [12, 13].
+For details on these methods, see Section˜B.2 and Appendix˜E.
+System 1: Meta-Learned Transformer Decomposer.
+System 1 is a decoder-only Transformer $\mathcal{T}_{\theta}$ trained not to solve any single task, but to perform a generic, single step of compositional decomposition.
+We meta-learn $\mathcal{T}_{\theta}$ on an endless stream of randomly generated grammars.
+Each training sample consists of a grammar definition, an input sequence, and the target output after one decomposition step at the highest-priority location.
+This forces the model to learn the abstract *process* of rule application rather than memorizing solutions for a specific task.
+See Section˜A.3 and Section˜A.4 for a full description of the training procedure and model vocabulary.
+Inference as a Processing Account.
+At inference time, System 2 first extracts the task grammar $\mathcal{G}$.
+Then, for a given input command, the model enters a loop (Figure˜1B) that serves as a processing account of step-by-step deliberation.
+In each step, the current sequence and the grammar are fed to the Transformer $\mathcal{T}_{\theta}$.
+For example, given ‘walk right twice after turn left’ and a grammar where ‘twice’ has higher priority than ‘after’, the Transformer identifies ‘walk right twice’ and rewrites it to a placeholder token, say P1.
+The sequence becomes ‘P1 after turn left’.
+In the next iteration, it decomposes this simpler structure.
+This process repeats until only a sequence of primitive actions remains.
+This iterative refinement allows a small model with a fixed context window to solve problems of arbitrary compositional depth.
+The full inference algorithm is detailed in Algorithm˜1.
+3 Experiments and Behavioral Analysis
+We evaluate Mirage on the canonical SCAN benchmark [4], which tests systematic generalization to novel compositions and longer sequences.
+The System 1 Transformer (1.19M parameters) is trained task-agnostically on random grammars; for evaluation, the System 2 Schema Engine extracts the SCAN grammar from its training split, which is then provided to the frozen Transformer.
+3.1 Compositional Generalization Performance
+As shown in Table˜1, Mirage achieves near-perfect accuracy on all SCAN splits, including those designed to foil standard sequence-to-sequence models.
+A vanilla Transformer trained directly on SCAN’s simple split masters it but completely fails on the generalization splits, indicating it learns superficial correlations rather than abstract rules.
+Providing the schema library as a simple prompt (‘Transformer+SC_Library’) does not help, confirming that a specialized architecture and training objective are required to properly utilize schematic knowledge.
+Table 1: SCAN accuracy (mean  $\pm$ SEM over 4 runs). Baselines are re-trained on each split’s training set. Mirage uses a single, task-agnostically trained Transformer for all splits.
+SCAN splits
+Model Full Task Simple Length Add prim. ‘jump’ Template
+Transformer N/A 99.8599.85± $\pm\,0.00$ 13.5813.58± $\pm\,0.01$ 0.400.40± $\pm\,0.13$ 3.093.09± $\pm\,3.01$
+Transformer+SC_Library N/A 99.9199.91± $\pm\,0.11$ 15.8615.86± $\pm\,1.36$ 0.030.03± $\pm\,0.04$ 0.000.00± $\pm\,0.00$
+Mirage (ours) 99.5999.59± $\pm\,0.24$ 99.5099.50± $\pm\,0.30$ 99.3599.35± $\pm\,0.41$ 99.6599.65± $\pm\,0.20$ 99.5599.55± $\pm\,0.23$
+3.2 Processing Account: Ablation Studies
+To understand which components drive this behavior, we conducted several ablations.
+These studies confirm that Mirage’s performance is an emergent property of its cognitive architecture, not just one component.
+  1. 1.
+    Schema priorities are crucial for cognitive control.
+Removing the explicit priority ordering from the schema library causes accuracy to plummet from 99.6% to 71.9%.
+The model becomes unable to resolve ambiguous commands (e.g., ‘turn left twice and walk’), analogous to a failure of executive function in selecting the correct order of operations.
+  2. 2.
+    Iterative refinement enables deliberation.
+Replacing the iterative loop with a single-pass, end-to-end mapping fails to generalize.
+The step-by-step process is essential, allowing the model to focus on one sub-problem at a time, a form of deliberate, sequential reasoning that monolithic models lack.
+  3. 3.
+    Accurate schemas are necessary.
+Introducing minor errors into the extracted grammar (e.g., misidentifying a schema’s arguments) causes performance to collapse to nearly 0%.
+This highlights the model’s reliance on accurate declarative knowledge.
+Its reasoning algorithm is sound, but it cannot function with a corrupted world model, a behavior consistent with human reasoning [14].
+4 Related Work
+Research on compositionality spans auxiliary objectives, neuro-symbolic models, meta-learning, and analyses of Transformer mechanisms.
+Auxiliary supervision can encourage structural understanding in Transformers [15], while memory-augmented models like LANE [16] explore variable-slot reasoning.
+Neuro-symbolic systems such as the Compositional Program Generator [17] and the Neural-Symbolic Recursive Machine [18] highlight the benefits of grammar-based modularity, paralleling Mirage’s schema-based design.
+Meta-learning approaches have also demonstrated strong potential for systematic generalization.
+In particular, Lake and Baroni [19] show that a meta-learned neural network can achieve human-like compositional generalization, supporting the view that inductive biases learned from experience can yield flexible reasoning skills.
+Transformer-focused studies reveal both strengths and limits: while large models can sometimes achieve compositional generalization [20], their reasoning often emerges only through grokking [21] or limited mechanistic patterns [22], underscoring the need for specialized architectures.
+Finally, neuroscience perspectives propose that hippocampal replay supports compositional computation [23], and that HPC–PFC circuits represent an evolutionary solution to flexible reasoning [24].
+These ideas motivate Mirage’s dual-process design, bridging cognitive neuroscience and AI.
+5 Discussion and Conclusion
+Mirage offers a hybrid approach to compositional generalization, using a neural network for a general procedural task (decomposition) while relying on an explicit, interpretable symbolic layer for declarative knowledge and control.
+Unlike monolithic models that must implicitly learn reasoning algorithms through “grokking,” Mirage explicitly encodes a process of deliberate, iterative refinement.
+This dual-process framework offers a concrete processing account for systematic generalization.
+By instantiating cognitive theories of complementary learning systems, our model achieves state-of-the-art zero-shot performance on SCAN.
+We release the full implementation of Mirage on GitHub.
+Our main contributions as an interpretive work are:
+  • Processing Account: We show that systematicity can emerge from an architecture that separates procedural skill from declarative knowledge and applies them iteratively.
+  • Behavioral Account: The model’s success where standard Transformers fail demonstrates the behavioral advantage of a neuro-inspired, modular design.
+  • Developmental Account: Meta-learning provides a hypothesis for how agents can acquire robust, generalizable cognitive algorithms from varied experience.
+  • Interpretability Account: Because Mirage explicitly marks and resolves schemas step by step, it produces a transparent trace of intermediate problem states.
+This contrasts with black-box models that leap directly to an answer, making our framework inherently more interpretable and diagnostically useful.
+Connection to Cognitive Learning Systems
+Beyond enabling stepwise compositional inference, schema formation as utilized by Mirage provides a mechanism for rapid assimilation of new rules into structured knowledge.
+In neuroscience, once an associative schema is established, new information consistent with that schema is integrated and becomes neocortically supported unusually quickly [10, 11, 25].
+Recent work further links medial prefrontal schema representations with reinforcement-learning style updating and deployment for flexible behavior [9].
+Mirage’s prioritized schema selection echoes this PFC-mediated control over hippocampal content during retrieval and planning [11].
+In addition, when viewed through a continual-learning lens, schemas could act as structured inductive priors: previously consolidated structure may be used to constrain and accelerate learning on new tasks.
+Bayesian treatments formalize this as using posteriors from past tasks to define priors for future ones, while also highlighting practical limits when approximations are poor [26, 27].
+These observations motivate future work in the direction of probabilistic extension: inferring over candidate schema targets (and, when needed, over the library itself) to maintain sample-efficient generalization when direct schema identification is challenging.
+Relation to recurrent/recursive Transformers
+A growing line of work explores iterative neural reasoning with weight-tying over steps, truncated credit assignment, and, in some cases, adaptive halting, yielding recurrent or recursive Transformer loops [28, 29, 30, 31].
+Mirage lies squarely within this trend by retaining an iterative per-step neural update, yet it differs in where the control signal lives: rather than learning control implicitly, we externalize it as explicit schemas with priorities.
+This keeps the empirical benefits of looped refinement while producing stepwise, editable traces and an algorithmic-level account of the reasoning process, which are central to Cognitive Interpretability.
+In this sense, Mirage complements recurrent and recursive Transformer practice with a cognitively motivated control layer, clarifying how and why each step is chosen.
+Limitations and Future Work
+While successful on SCAN, the current framework relies on reliable extraction of a consistent grammar.
+Future work should explore robustness to noisy or incomplete schemas.
+Such issues could potentially be addressed by extending the extractor with probabilistic rule learning and/or the Neural Decomposer with active retrieval mechanisms.
+We also plan to apply this architecture to more complex reasoning tasks and use its inspectable states to test specific neuroscientific hypotheses about HPC–PFC interactions during planning.
+By building explicit, interpretable cognitive architectures, we can move beyond black-box evaluations and toward a mechanistic understanding of reasoning in neural systems.
+Appendix A Methodology details
+A.1 Formal definition of Transformer Inference
 Algorithm 1 Zero-Shot Inference on Single Input Sequence Example
@@
 5:for $d=k,k-1,\dots,1$ do
-6:  $y\leftarrow\mathcal{T}_{\theta}\bigl{(}x^{(d)}\bigr{)}$triangleright predict next-level decomposition
+6:  $y\leftarrow\mathcal{T}_{\theta}\bigl(x^{(d)}\bigr)$ $\triangleright$ predict next-level decomposition
 7:  while there is a schema token $s$ in $y$ do
 8:   let $(\sigma_{s},\,a_{1},\dots,a_{m})$ be the schema name and its $m$ argument tokens in $y$
-9:   $p\leftarrow\texttt{ApplySchema}(\sigma_{s},\,a_{1},\dots,a_{m})$triangleright collapse into one primitive
+9:   $p\leftarrow\texttt{ApplySchema}(\sigma_{s},\,a_{1},\dots,a_{m})$ $\triangleright$ collapse into one primitive
 10:   $\lambda\bigl[\sigma_{s}(a_{1},\dots,a_{m})\bigr]\leftarrow p$
 11:   replace the subsequence $\langle\sigma_{s},a_{1},\dots,a_{m}\rangle$ in $y$ with $p$
-12:  $x^{(d-1)}\leftarrow y$triangleright one less level of composition
+12:  $x^{(d-1)}\leftarrow y$ $\triangleright$ one less level of composition
 13:return $\mathcal{S}_{\mathrm{final}}\leftarrow x^{(0)},\;\lambda$
-After receiving the output from algorithm 1, we simply resolve or unwind the complete value for $S_{final}$, replacing any tokens in the sequence that match keys in $\lambda$ with their corresponding value in $\lambda$.
-The number of model iterations required to completely resolve a schema is equal to the depth of the composition tree, where the root node sits at depth 0.
-Consider an input sequence of length $N$, with a max number of arguments per schema of $A$ (in SCAN, $A=2$).
-Then, the depth of the tree is $O(\log_{A}N)$, and as such, the number of decomposition or reasoning steps is generally logarithmic in sequence length.
-This allows processing extremely large compositional sequences in a small number of reasoning steps.
-2.5 System 2 details: complementary schema extractors
-Schema extraction is treated as a modular component whose internal strategy can be varied without altering the rest of Mirage.
-We currently provide two contrasting implementations: a symbolic search procedure and a learnable graph-based model built atop a recent computational model of the hippocampus George et al.
-(2021).
-Their different strategies make it unlikely that both will fail on the same input, and either one can be replaced as improved algorithms or neuroscientific insights become available.
-Extractor Option One: CSCG-inspired extractor
-Following the neuroscientific inspiration of the two-component model, we utilize Clone-Structured Causal Graph (CSCG) George et al.
-(2021), effective in modeling cognitive maps in mice Sun et al.
-(2025).
-Adapted from Cloned Hidden Markov Models (CHMM) Dedieu et al.
-(2019), CSCGs efficiently represent complex sequences via a discrete latent state model.
-CSCGs further constrain each latent state to deterministically emit a single token, effectively ‘cloning’ each token into a large number of hidden states.
-Critically, CSCGs can rebind these emission matrices to generalize to patterns analogous to those they were trained on, allowing them to learn simple algorithms Swaminathan et al.
-(2023).
-Our approach works by extending this rebinding capability.
-The original rebinding mechanism triggers upon encountering unexpected tokens and depends on stable “anchor” tokens that are rarely rebound.
-We extend this mechanism to bidirectional contexts where novel primitives might precede anchors.
-Thus, we introduce an adapted rebinding algorithm to explicitly detect anchors and rebinding slots within such contexts, defined formally below.
-[…]
-1.
-Episode decoding.
-We train a CHMM on the raw token stream.
-Viterbi decoding yields a latent state trace for each demonstration, forming episodes $e=\langle\mathbf{x},\mathbf{y},\mathbf{z}\rangle$ where $\mathbf{z}$ is the state sequence.
-These episodes serve as structured surrogates of hippocampal encodings.
-2.
-Schema discovery and pruning.
-Candidate schemas are generated by aligning episode pairs $(e_{i},e_{j})$ using the longest common subsequence (LCS) under a constraint enforcing consistent variable bindings.
-A candidate is accepted if it validates on all supporting episodes via forward simulation and is supported by at least min_support examples, filtering out ambiguous or redundant rules.
-Surviving schemas are deduplicated using a keyword–coverage heuristic and pruned iteratively if compositionally subsumed by higher-support schemas.
-We visualize the set of extracted schemas in Appendix Figure 4.
+A.2 Formal definition of CSCG-based System 2 Algorithm
 Algorithm 2 Extraction of compositional schemas
@@
 4:for all $(e_{i},e_{j})\in\binom{\mathcal{E}}{2}$ do
-5:  $\tau\leftarrow\textsc{Align}(e_{i},e_{j})$triangleright injective LCS with variables
-6:  if $\tau\not=\bot$ then
+5:  $\tau\leftarrow\textsc{Align}(e_{i},e_{j})$ $\triangleright$ injective LCS with variables
+6:  if $\tau\not$$\bot$ then
 7:   $\text{supp}\leftarrow\{e\in\mathcal{E}\mid\textsc{Validate}(\tau,e)\}$
@@
 11:return $\mathcal{S}_{\text{final}}$
-Schemas are applied via a greedy, pass-based interpreter that replaces the longest matching input spans until convergence.
-A derivation is successful if all placeholders are resolved.
-In compositional cases like SCAN, however, overlapping operator-style schemas (e.g., turn right thrice) may conflict.
-To resolve this, we induce a precedence relation $\succ$ over $\mathcal{S}_{\text{final}}$.
-Inspired by CSCG-style rebinding, we use comparative demonstrations split by <SEP> and utilize one-step resolution.
-If schema $s_{i}$ fires before $s_{j}$, we record $s_{i}\succ s_{j}$.
-Aggregating such wins produces a matrix $C\in\mathbb{N}^{|\mathcal{S}_{\text{final}}|\times|\mathcal{S}_{\text{final}}|}$, whose Copeland score
-$\text{score}(s_{i})=\sum\nolimits_{j\neq i}\bigl{[}\![\,C_{ij}>0\,]\!-\![\,C_{ji}>0\,]\bigr{]}$
-defines a total preorder. Full details are in Appendix B.1.
-With precedence learning in place our system (i) discovers variable-binding rewrite rules, (ii) prunes them for compositional minimality, and (iii) orders overlapping operator schemas so that the following Transformer model can utilize deterministic schema information on previously unseen commands.
-Extractor option two: enumerative rule miner.
-To complement the CSCG path we include a lightweight symbolic “rule-miner” inspired by enumerative program-synthesis methods Udupa et al.
-(2013).
-Starting from the demonstrations alone, it treats each input–output pair as an episodic memory and enumerates simple string-rewrite templates (span-to-token replacements, span re-ordering, and wrapper insertions).
-Templates are accepted only if, when added to the current library, they repair all of their matches and raise corpus-level exact accuracy; accepted rules are then replayed to expose new residual errors and the loop repeats until coverage stabilizes.
-The miner also records “fires-before” relations between overlapping rules and returns a topologically-sorted precedence schedule together with the final library of primitives, modifiers, and schemas.
-The resulting grammar is directly fed into the Schema Engine and gives performance indistinguishable from the CSCG variant, underscoring System 2’s modularity.
-3 Results
-3.1 MIRAGE Performance on SCAN vs. Baselines
-We evaluate Mirage and baselines on the canonical SCAN benchmark, following the exact data splits of Lake and Baroni (2018).
-Unless stated otherwise, all numbers are means $\pm$ SEM over 4 independent runs.
-Training, hardware, and hyper-parameter details mirror those in §2.
-Table 1 summarizes our main results.
-Mirage achieves 99.59 $\pm$ 0.24% accuracy on the full task while trained in a task-agnostic manner on randomly generated grammars rather than SCAN-specific examples, and reaches near-perfect performance on every split.
-Perhaps even more importantly, a single trained Mirage is generalizable to any compositional grammar (given its specification) with appropriate vocabulary size for the trained model (number of modifier tokens, primitives, and arguments per schema must be less than or equal to the corresponding parameters used to train the model).
-Table 1: SCAN accuracy (mean  $\pm$ SEM over 4 runs). † Baseline is re-trained on each split’s train set;
-SCAN splits
-Model Full Simple Length prim_jump temp_or
-[…]
-[…]
-[…]
-In contrast, transformers trained directly on SCAN solve the normal split effectively, but fail to generalize across length and new templates in SCAN.
-This demonstrates significant overfitting, failing to learn any generalizable notion of composition or compositional reasoning.
-The Transformer+SC_Library baseline represents a standard transformer trained to evaluate SCAN, with the addition of the SCAN schema library in-context (i.e.
-pre-pend the SCAN schema library tokens used for inference evaluation of MIRAGE to the training sequences of the transformer).
-We compare against a simple baseline by prepending the schema library directly to the input sequence to test whether the Transformer can interpret it without explicit guidance.
-Interestingly, as shown in Table 1, the opposite holds true: the additional context reduces performance on token-based splits and leads to only minor gains on the length split.
-This suggests that the model cannot effectively leverage schema information without a dedicated training objective, and length split improvements appear to merely stem from the increased input size rather than meaningful schema usage.
-3.2 Ablation Analysis Study
-To identify which design choices are essential for MIRAGE, we ran four ablations: (1) removing priority tokens, (2) disabling iterative refinement and forcing single-pass decomposition, (3) training on sequences with unbounded composition depth, and (4) evaluating with imperfectly extracted grammars.
-Each change substantially degraded accuracy, underscoring the importance of explicit precedence cues, step-wise decomposition, bounded-depth meta-training, and a correct schema specification.
-We have the following key observations:
-  1. 1.
-    Priority Tokens are Critical: Eliminating explicit priority scheduling reduces full-task accuracy from 99.59 ± 0.24% to 71.92 ± 0.72%.
-Although the model still solves a majority of commands, the 28-point drop reveals that it often chooses the wrong schema when several overlap.
-In practice, the transformer falls back on brittle positional cues and succeeds only when those cues coincide with the intended hierarchy.
-Placing schemas in a fixed order without dedicated PRIORITY_ tokens proved insufficient during development, confirming that the network needs explicit embeddings to internalize precedence.
-Prioritization therefore remains the hardest part of single-step decomposition; giving the model its own priority-token embedding space was crucial for reliable generalization.
-  2. 2.
-    Iterative Refinement Process Facilitates Generalization: We experimented with parameterizing complete single-shot decomposition prior to switching to single-step decomposition with the iterative refinement algorithm described above.
-In our experiments, we failed to train transformers to systematically learn or generalize complete single-shot decomposition algorithms even when trained on $k$-deep composition example sequences (2-deep train sequences also failed).
-The iterative refinement algorithm allows transformer models to learn more generalizable rules focused solely on prioritized schema identification, as opposed to full decomposition.
-In this sense, the iterative refinement algorithm seems important for facilitating generalized inference.
-  3. 3.
-    Training with Unlimited Depth Composition Sequences Fails: While we cannot present specific evidence as to why this might be the case, experiments repeatedly revealed that models (especially, larger models) trained on sequences including $k$-deep composition (limited only by sequence context length) tended to perform worse on SCAN, reaching a standard peak performance of less than 50%.
-Training on 2-deep composition sequences ultimately proved more generalizable even when performing inference on k-deep composition sequences, like in SCAN.
-This may be the result of models trained on deeper sequences attempting to model composition more closely, which ultimately becomes detrimental, as general patterns are not observed.
-A more focused learning paradigm on prioritized pattern matching only proves more generalizable and effective.
-  4. 4.
-    Correct Grammar Libraries are Necessary: Introducing stochastic boundary noise, that is, random shifts of one or two positions in num_args_before or num_args_after for every schema, drops full-task accuracy from 99.59 ± 0.24% to 0.065 ± 0.021%.
-Across four random seeds, the accuracy never exceeded 0.11 percent.
-This sharp decline shows that accurate schema extraction is essential; even small mistakes in the grammar cause the model’s compositional reasoning to break down.
-3.3 Experimental Takeaways
-Our findings support the hypothesis that coupling an explicit schema extraction system with a fast neural processor is necessary and (empirically) sufficient for systematic compositionality.
-Our contributions may be summarized through the following key points.
-(1) A HPC-PFC-inspired model consisting of corresponding components can effectively solve compositional generalization tasks zero-shot, when given a complete grammar specification.
-(2) The HPC-PFC complex, System 2, may be modeled with a wide variety of potential schema extraction methods.
-We present two methods based on program synthesis and neuro-inspired CSCG models.
-(3) Meta-learning transformers on an infinite random grammar stream with only 2-level deep composition training examples is sufficient to perform generalized single-step decomposition on arbitrary grammars in k-deep composition scenarios.
-Via an iterative refinement process, a trained model $\mathcal{T}_{\theta}$ can effectively decompose length $N$ compositional sequences in $O(\log N)$ reasoning steps.
-(4) We provide ablation results demonstrating the importance of the 2-system design and other components integrated into our core system.
-We believe that these contributions, specifically the overall framework of our system design and the use of transformers in an iterative reasoning/refine process, present intriguing areas of research moving forward.
-We believe that developing principled neuro-inspired AI systems represents perhaps the most tangible pursuit of the type of systematic generalization to new contexts and scenarios that humans excel at and even massive scale AI has not yet seemed to achieve.
-4 Related Work
-Adjacent Work on Compositionality
-Several research directions have tackled compositional generation, spanning auxiliary objectives, neuro-symbolic models, prompting strategies, and mechanistic Transformer analyses.
-Jiang and Bansal (2021) use auxiliary supervision to improve compositional generalization in Transformers, by introducing sequence prediction tasks that encourage structural understanding, finding that less contextualized representations improve SCAN performance.
-Another strategy, LANE Liu et al.
-(2020), learns analytical expressions and leverages memory augmentation inspired by variable-slot reasoning from cognitive science.
-Though related to our extractor component, LANE is non-Transformer-based and follows a distinct training paradigm, yet still achieves strong SCAN results.
-Neuro-symbolic methods combine neural and symbolic reasoning for compositionality.
-The Compositional Program Generator (CPG) Klinger et al.
-(2024) integrates grammar-based modularity and abstraction, aligning closely with Mirage’s schema-based extraction strategy.
-Similarly, the Neural-Symbolic Recursive Machine (NSR) Li et al.
-(2024) learns syntax and semantics jointly via a Grounded Symbol System, using tree structures reminiscent of our Transformer-based neural processor.
-Compositionality in Deep Learning Structures
-Finally, understanding the capabilities and limitations of existing architectures, particularly Transformers, is essential for developing more effective models.
-Russin et al.
-(2024) survey recent work on compositionality in deep neural networks (DNNs), exploring how DNNs, particularly large language models (LLMs), can achieve compositional generalization to a specific degree through architectural inductive biases and metalearning.
-Wang et al.
-(2024) investigate Transformers’ ability to perform implicit reasoning, finding that they can learn these skills only through "grokking", but struggle with systematic generalization in compositional tasks.
-Brinkmann et al.
-(2024) perform a mechanistic analysis of a Transformer trained on a synthetic reasoning task, identifying interpretable mechanisms like backward chaining.
-Overall, these works highlight the challenges in achieving robust compositional reasoning with standard transformers, further motivating the need for specialized architectures like Mirage.
-Neuroscience Inspiration
-We further take some to reflect upon work in neuroscience and cognition to discuss the inspiration Mirage was built upon.
-Kurth-Nelson et al.
-(2022) propose that hippocampal replay implements compositional computation by assembling entities into relationally-bound structures, aligning with Mirage’s schema application and transformation mechanisms.
-Their hypothesis strengthens the neuroscientific grounding of Mirage by linking it to the compositional role of replay in the brain.
-The HPC-PFC circuits, conserved across mammalian evolution Murray et al.
-(2017), represent nature’s time-tested solution to compositional reasoning, offering a bridge between neuroscience and AI that may illuminate new paths toward more human-like intelligence.
-5 Conclusion
-We introduced Mirage, a neuroscience-inspired dual-process framework that explicitly separates fast pattern matching from deliberate schema manipulation.
-Meta-learning a single-step Transformer decomposer over a stream of random grammars, combined with a priority-aware schema manager, yields state-of-the-art zero-shot compositional generalization, given a complete grammar: Mirage attains 99.6% accuracy on the full SCAN task and performs similarly on the other splits, without exposing SCAN training data to the Transformer.
-Beyond outperforming strong baselines, our ablation study confirms that (i) priority scheduling and (ii) iterative refinement are each critical: Removing either component collapses systematicity.
-Looking more broadly at the reasoning landscape today, dominant approaches operate by training large homogeneous neural networks to perform step-by-step inference Guo et al.
-(2025); Muennighoff et al.
-(2025); OpenAI et al.
-(2024).
-Our model should be seen as an alternative route.
-We learn a small step-by-step reasoner, and augment it with an external module inspired by the modular architecture of the brain.
-Our model also sits within the paradigm of separating knowledge from inference, which is a staple of classic logical reasoning.
-Looking forward we hope that the judicious application of modularity, particularly inspired by brain architecture, could prove a promising complementary route for learning to reason.
-Limitations.
-We initially consider a radically general knowledge representation—term rewriting systems—but then quickly restrict it to make the model tractable.
-We are optimistic however that taking Turing-complete representations as a starting point could prove valuable in the future.
-As our modeling goals are primarily inspired by neuroscience, we also take a noncommittal stance regarding the optimal algorithm for learning or extracting schemas, and provide two distinct algorithms.
-For CSCG extraction we further assume existence of atomic demonstrations.
-Last, we only evaluate on SCAN, because it is a canonical check of compositionalty Lake and Baroni (2023), but we hope that by moving toward increasingly expressive computational formalisms for schemas, our architecture could find further application.
-Future Directions and Broader Impacts.
-We plan to (i) embed Mirage as a plug-in reasoner/planner for LLMs, while aiming to enhance their abilities to more robustly build and use world models; and (ii) examine its internal states to test neuroscientific hypotheses about HPC–PFC interactions.
-Broader impact: by exposing an inspectable schema layer, Mirage can make neural reasoning more transparent and auditable, however, deployments, particularly in safety-critical settings, should include clear usage policies and routine monitoring.
-Appendix
-Appendix A Methodology details
-A.1 Transformer Model Vocabulary Specifications for Meta-Learning
+A.3 Transformer Model Vocabulary Specifications for Meta-Learning
 The vocabulary of our transformer $\mathcal{T}_{\theta}$ is defined with the following components.
@@
     Administrative Tokens: EOS, SEP, SC_DEF, SC_PRI, SC_SEP, LP_SEP , PAD. These tokens are for formatting grammars in-context with input sequences generated by them.
-A.2 Training Algorithms
+A.4 Training Algorithms
 During the transformer training process, new random grammars are generated at regular intervals.
@@
 Algorithm 3 Learning operator precedence from one-step demonstrations
-[…]
+1:fixed schema set $\mathcal{S}$, demonstration episodes $\mathcal{E}$ of the form $\langle\text{cmd}\rangle\!<\text{SEP}>\!\langle\text{1-step}\rangle$
 […]
@@
 4:  $(\mathbf{x},\mathbf{y})\leftarrow$ split $e$ at <SEP>
-5:  $P\leftarrow\{\,s\in\mathcal{S}\mid\text{pattern}(s)\subset\mathbf{x}\,\}$triangleright schemas present in the input command
-6:  $F\leftarrow\{\,s\in P\mid\text{apply}(s,\mathbf{x})=\mathbf{y}\,\}$triangleright schemas that fired in the first step
+5:  $P\leftarrow\{\,s\in\mathcal{S}\mid\text{pattern}(s)\subset\mathbf{x}\,\}$ $\triangleright$ schemas present in the input command
+6:  $F\leftarrow\{\,s\in P\mid\text{apply}(s,\mathbf{x})=\mathbf{y}\,\}$ $\triangleright$ schemas that fired in the first step
 7:  for all $s_{w}\in F$ do
@@
 10:for all $s_{i}\in\mathcal{S}$ do
-11:  $\text{score}(s_{i})\leftarrow\sum_{j\neq i}\bigl{(}[\![C_{i,j}>0]\!]-[\![C_{j,i}>0]\!]\bigr{)}$triangleright Copeland score
+11:  $\text{score}(s_{i})\leftarrow\sum_{j\neq i}\bigl([\![C_{i,j}>0]\!]-[\![C_{j,i}>0]\!]\bigr)$ $\triangleright$ Copeland score
 12:return $\{\text{score}(s_{i})\}_{s_{i}\in\mathcal{S}}$
 B.2 Enumerative Rule-Miner Details
-This appendix expands on the enumerative rule miner introduced in §2.5.
+This appendix expands on the enumerative rule miner introduced in Section˜2.
 B.2.1 Goal
@@
   1. (1)
-[…]
+    Template generation Enumerate candidate LHS–RHS templates in order of description length via three language-agnostic edit primitives: (i) span $\to\!$ token replacements, (ii) span wrappers inserting a control token, and (iii) span splicing / re-ordering.
   2. (2)
@@
 Figure 2: We apply a Clone-Structured Causal Graph with 100 clones directly on a concatenated subset of SCAN sequences and visualize the resulting model as a simple directed graph.
-In line with showcasing the inability of pure Transformers to solve different SCAN splits on its own in section 3, we explored directly applying CSCG on concatenated SCAN sequences.
+In line with showcasing the inability of pure Transformers to solve different SCAN splits on its own in Section˜3, we explored directly applying CSCG on concatenated SCAN sequences.
 However, due to innate traits of the model architecture, SCAN, or other compositional tasks, quickly exhaust the structural clone bottlenecks.